OpenStudy (anonymous):

Find the two-digit numeral whose units' digit is 5 more than its tens' digit and whose value equals 3 times the sum of these digits.

OpenStudy (anonymous):

Let the ten's be x, then its units are x+5 and the sum of digits is x+x+5=2x+5 and the value of a number will be:10x+x+5=11x+5. Since its value equals 3 times the sum of the digits we write the equation: \[11x + 5 = 3\times(2x + 5)\] \[5x = 10\] x = 2 The units are 2+5=7. Answer: The two digit number is 27.

OpenStudy (anonymous):

Thanks a million!!! I was wondering is there a way for me to come back for your help later? I have to go to church.